evaluate ∫ 2 cos^2 x- cos 2x / cos^2 x dx

evaluate
∫ 2 cos^2 x- cos 2x / cos^2 x  dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \frac{2cos^{2}x - cos2x}{cos^{2}x}dx
cos2x = 2cos^{2}x - 1
2cos^{2}x - cos2x = 1
I = \int \frac{1}{cos^{2}x}dx
I = \int sec^{2}x.dx
I = tanx + constant